OpenStudy (winsteria):

A recipe for potting soil that makes enough to fill 5 bags calls for 3 (and) 3/4 lb of compost. Michael wants to modify the recipe so that it fills 14 bags. How many more pounds of compost will Michael need for the modified recipe than the original recipe calls for? A. 11 (and) 1/4 B. 10 (and) 1/2 C. 7 (and) 1/4 D. 6 (and) 3/4 CAN SOMEONE PLEASE HELP ME

OpenStudy (tkhunny):

Find a unit rate: 5 bags / 3.75 lb = 4/3 bag/lb OR Use a proportion: 5/3.75 = 14/What?

OpenStudy (winsteria):

so WHATS THE ANSWER??? D:

OpenStudy (winsteria):

what letter???

OpenStudy (tkhunny):

The correct solution is the answer. You tell me.

OpenStudy (tkhunny):

I don't actually care. I didn't look at the answers. My job is to help YOU solve the problem.

OpenStudy (winsteria):

Oh okay well... uh I think that it's 6 (and) 3/4's but I don't really know D:

OpenStudy (winsteria):

CAN SOMEONE HELP ME PLEASEEEEEEE

OpenStudy (tkhunny):

How did you get that?

OpenStudy (winsteria):

I tried to divide it and i'm mostly just trying to guess because I need the answer fast so D:

OpenStudy (tkhunny):

You tried to divide what? Please make an attempt to communicate useful information. Look really hard at that proportion.

OpenStudy (winsteria):

I'm pretty sure that it's A know which is 11 (and) 1/4 and I don't really know how to explain how I got the answer but that's just what I think it is.

OpenStudy (tkhunny):

Good thinking, but poor calculating. You did NOT look at that proportion. It might be easier the ether way. 3.75/5 = What/14